OpenStudy (anonymous):

What is the enthalpy change when 13.8 grams of H2O changes from a liquid to a solid at zero degrees Celsius? ΔHcondensation = −40.65 kJ/mol ΔHfreezing = −6.03 kJ/mol −4.62 kJ −83.2 kJ −7.87 kJ −31.3 kJ

OpenStudy (jfraser):

how many moels of water is in 13.8g of H2O?

OpenStudy (anonymous):

there is 1.22 moles

OpenStudy (jfraser):

no. 1mol of H2O is 18g. You've only got 13.8g. Flip the fraction upside down

OpenStudy (anonymous):

okay so the answer would −7.87 kJ ?

OpenStudy (anonymous):

or am i wrong?

OpenStudy (jfraser):

your mole conversion should look like\[13.8g H_2O * (\frac{1mol H_2O}{18gH_2O}) = 0.767 moles H_2O\]

OpenStudy (anonymous):

so is my answer wrong?

OpenStudy (jfraser):

mole of water "costs" 40.65kJ of energy to freeze. You've got only 76.7% of 1 mole. The heat of fusion for .767mol of H2O will be \[0.767mol H_2O * (\frac{40.65kJ}{1mol H_2O}) = 31.17kJ\]

OpenStudy (mos1635):

40.65 or 6.03 ????

OpenStudy (anonymous):

Oh okay . −31.3 kJ so this would be the final result?

OpenStudy (jfraser):

@mos1635 is correct, replace the 40.65kJ with 6.03kJ. I missed the condensation vs freezing.

OpenStudy (mos1635):

-4.623\[\approx\] -4.62